Solve: log(4x+1)-log(2x+9)=0 for x.
Answer by lwsshak3(11628)

Solve: log(4x+1)-log(2x+9)=0 for x.
place under a single log
log[(4x+1)/(2x+9)]=0
convert to exponential form: base(10) raised to log of number(0)=number[(4x+1)/(2x+9)]
10^0=[(4x+1)/(2x+9)]=1
(4x+1)=(2x+9)
2x=8
x=4